A twenty-year-old decides to save Rs.1,000 per month into a retirement account that will
mature when she will turn 65. How much will she accumulate assuming that the she gets
a return @12% per annum compounded monthly? How much will a 40-year-old get if he
makes a similar investment? How much more does he need to invest to get the same
amount?
1
Expert's answer
2020-12-17T07:43:20-0500

in 65 years:

"1000*\frac{(1+0.12/12)*((1+0.12/12)^{(65-12)*12}-1}{1.01-1}=56,485,795.36"

40-year-old:

"1000*\frac{(1+0.12/12)*((1+0.12/12)^{(65-40)*12}-1}{1.01-1}=1,897,635.09" - will get

needs to invest per month:

"56,485,795.36-1,897,635.09=54,588,160.27"

"54,588,160.27/\frac{(1+0.12/12)*((1+0.12/12)^{(65-40)*12}-1}{1.01-1}=29,054.08"